OpenStudy (fvf1977):

Solve the equation exactly in radians: cos^2x=1/2sin(2x), all x

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \cos^2x=\frac{1 }{2\sin(2x)} }\) like this?

OpenStudy (fvf1977):

sin(2x) is in the numerator

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \cos^2x=\frac{\sin(2x) }{2} }\) right?

OpenStudy (fvf1977):

\[\cos ^{2}x=(1/2)\sin(2x)\]

OpenStudy (solomonzelman):

yes, I got it then, it is the same exact thing

OpenStudy (solomonzelman):

There is a rule that: \(\large\color{brown}{ \displaystyle \sin(2x)=2\sin(x)\cos(x) }\) first, rewrite it like this

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \cos^2x=\frac{2\sin x \cos x }{2} }\)

OpenStudy (solomonzelman):

the 2's cancel, and you get: \(\large\color{black}{ \displaystyle \cos^2x=\sin x \cos x }\)

OpenStudy (solomonzelman):

can you finish from here?

OpenStudy (solomonzelman):

(subtract sin(x)cos(x) from both sides and factor)

OpenStudy (fvf1977):

I understand so far & no:(

OpenStudy (fvf1977):

I can factor out a cosx right?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \cos^2x=\sin x \cos x }\) \(\large\color{black}{ \displaystyle \cos^2x-\sin x \cos x =0 }\) yes, factor out of cos(x)

OpenStudy (fvf1977):

so i should get cosX(cosX-sinX)=0

OpenStudy (solomonzelman):

yes, correct

OpenStudy (solomonzelman):

then, either cos(x)=0 OR cos(x)-sin(x)=0

OpenStudy (solomonzelman):

solve each of them

OpenStudy (fvf1977):

for cos(x)=0 i got \[\pi/2\] & \[3\pi/2\]

OpenStudy (solomonzelman):

wait, your interval is then supposed to be \([0,2\pi]\) ?

OpenStudy (solomonzelman):

are you looking for solutions over this (or some other) interval, or a general solution?

OpenStudy (fvf1977):

it says "all X"

OpenStudy (solomonzelman):

then you need all possible answers

OpenStudy (solomonzelman):

For \(\cos(x)=0\) , your solutions are: \(\pi/2\), \(\pi/2~~+\pi\), \(\pi/2~~+2\pi\), \(\pi/2~~+3\pi\), and so on.... right?

OpenStudy (solomonzelman):

and so would be true if you subtract Pi's instead of adding them

OpenStudy (solomonzelman):

So, \(\LARGE x={\rm n}\cdot\pi+\frac{1}{2};~~~\left\{n\in {\bf Z}\right\}\)

OpenStudy (solomonzelman):

the last { } tell you that "n" can be any integer

OpenStudy (fvf1977):

okay and cosX-sinX= 0, how do you do that one

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \cos(x)-\sin(x)=0 }\) \(\large\color{black}{ \displaystyle \cos(x)=\sin(x) }\) do you remember the angle where sine and cosine functions are equal ?

OpenStudy (fvf1977):

pi/4?

OpenStudy (solomonzelman):

yup, there you go

OpenStudy (solomonzelman):

can you give me some other examples?

OpenStudy (fvf1977):

5pi/4?

OpenStudy (solomonzelman):

yes, (you are adding pi), very good

OpenStudy (solomonzelman):

or, you can have -3pi/4 , can you?

OpenStudy (solomonzelman):

So you can go on subtracting \(\pi\) or adding \(\pi\), right ?

OpenStudy (fvf1977):

Yes but how do you know the equation to write? it always confuses me like the solution for cosX=0 and you have multiple solutions. if that makes sense

OpenStudy (solomonzelman):

that is adding \(n\cdot \pi\) (where when n is a negative integer you are subtracting)

OpenStudy (solomonzelman):

you have multiply solutions because you can go around many times (if you imagine a unit circle)

OpenStudy (fvf1977):

Our teacher uses 'K' and just says which K is any real #

OpenStudy (solomonzelman):

these are the "coterminal angles"

OpenStudy (solomonzelman):

K instead of n, then we can write K

OpenStudy (solomonzelman):

capital or lowercase k ?

OpenStudy (fvf1977):

Capital

OpenStudy (solomonzelman):

Okay, please try to write the solution set for \(\large\color{black}{ \displaystyle \cos(x)=\sin(x) }\)

OpenStudy (solomonzelman):

(I will correct if anything)

OpenStudy (solomonzelman):

well, what is the pattern, can you tell me?

OpenStudy (fvf1977):

adding pi?

OpenStudy (solomonzelman):

yes, but you are adding any number of pi, you are adding 1pi , 2pi, 3pi, and so on..... OR, you are subtracting 1pi, 2pi, 3pi, (which is, adding -1pi, -2pi, -3pi ... and on) \(\large\color{black}{ \displaystyle \frac{ \pi }{4} }\) this is our starting point, and then you are adding these negative or positive pi's \(\large\color{black}{ \displaystyle \frac{ \pi }{4} +K\pi }\) (K can be negative, because it is an integer; this pattern perfectly satisfies the solution to the equation)

OpenStudy (solomonzelman):

And then add a notation that K in an integer \(\large\color{black}{ \displaystyle \frac{ \pi }{4} +K\pi;~~n\in{\bf Z} }\)

OpenStudy (solomonzelman):

If anything seems not very clear, please ask....

OpenStudy (fvf1977):

Ok understood and the first one (cosX=0) how does the equation look with K?

OpenStudy (solomonzelman):

I thought I have already given it, ... I will go through that one anyway

OpenStudy (solomonzelman):

your starting point is Pi/2 then you add or subtract K\(\pi\) from that (where K is an integer)

OpenStudy (solomonzelman):

without looking back into the thread, can you write the pattern for me please?

OpenStudy (fvf1977):

\[\pi/2+K \pi \]?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

where K is an integer

OpenStudy (solomonzelman):